# Sylber: Syllabic Embedding Representation of Speech from Raw Audio

### v1

@@ Abstract
Syllables are compositional units of spoken language that play a crucial role in human speech perception and production. However, current neural speech representations lack structure, resulting in dense token sequences that are costly to process. To bridge this gap, we propose a new model, Sylber, that produces speech representations with clean and robust syllabic structure. Specifically, we propose a self-supervised model that regresses features on syllabic segments distilled from a teacher model which is an exponential moving average of the model in training. This results in a highly structured representation of speech features, offering three key benefits: 1) a fast, linear-time syllable segmentation algorithm, 2) efficient syllabic tokenization with an average of 4.27 tokens per second, and 3) syllabic units better suited for lexical and syntactic understanding. We also train token-to-speech generative models with our syllabic units and show that fully intelligible speech can be reconstructed from these tokens. Lastly, we observe that categorical perception, a linguistic phenomenon of speech perception, emerges naturally in our model, making the embedding space more categorical and sparse than previous self-supervised learning approaches. Together, we present a novel self-supervised approach for representing speech as syllables, with significant potential for efficient speech tokenization and spoken language modeling.
@@ 1 Introduction
A major bottleneck of the inefficiency in modeling spoken language is a lack of structure in current neural speech representations. Unlike text, there is no clear delimiter nor orthographic symbol in speech audio, which are crucial in efficient and scalable processing as evidenced in the text domain. However, human speech perception is structured as being segmented (Greenberg, 1998; Oganian & Chang, 2019; Gong et al., 2023) and categorical (Liberman et al., 1957; Pisoni, 1973; Pisoni & Lazarus, 1974). We argue that the machine representation of speech should resemble these cognitive structures to allow similar efficiency as text processing. A natural segmented structure of speech is a syllable, which organizes speech sounds in time (MacNeilage, 1998; Greenberg, 1998), and ideally, the embedding of a syllable should represent contents in a categorical way to avoid redundancy prevailing in current SSL-based tokens.
To this end, we propose self-segmentation distillation, a novel SSL framework that induces clean and robust syllabic structures in speech representations. Specifically, we build on top of a previous self-supervised syllable learning model, SDHuBERT (Cho et al., 2024b), and iteratively refine the syllabic segments that naturally arise from the model. Unlike the original model, which induces syllable structure as a byproduct of sentence-level SSL, we directly impose syllabic structures by regressing features against unsupervised syllable segments extracted from a teacher model which is a moving average of the training model. We call the resulting model Sylber (Syllabic embedding representation). [^1] The code is available here: https://github.com/Berkeley-Speech-Group/sylber
The features from Sylber exhibit salient syllabic structure — showing a flat, consistent output within each segment and distinctive from other syllables (Figure 2, right). This enables a fast, linear time algorithm for segmenting these features. Moreover, this allows more accurate boundary detection and clustering that is more coherent with ground truth syllables than previous approaches. Syllabic tokens quantized from Sylber features show significantly lower frequency at an average of 4.27 token/second, and can be used to synthesize fully intelligible speech. [^2] Audio samples are at https://berkeley-speech-group.github.io/sylber and in supplementary materials. Furthermore, unit LMs based on syllabic tokens outperform the baselines with a similar resource setting, in learning lexicons and syntax.
@@
  • We propose self-segmentation distillation, a novel SSL framework that imposes salient and robust syllabic structure in speech representation.
  • The resulting model, Sylber, outperforms previous approaches in syllable detection and discovery with a segmentation algorithm with $O(n)$ time complexity.
  • We use this model to build a syllable-level speech tokenization scheme that has significantly lower sampling rate as 4.27 Tok/s on average, 6-7 times improvement over previous HuBERT tokens.
  • We demonstrate that fully intelligible speech can be reconstructed from syllabic tokens, and that these units are better suited for spoken language understanding.
@@ 2 Related Work
Self-supervised learning in speech Self-supervised learning (SSL) has been leveraged in speech to learn representations from large, unlabeled speech corpuses (Hsu et al., 2021; Baevski et al., 2020; Chen et al., 2022; Chung et al., 2021; Mohamed et al., 2022). Notably, HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022) are pretrained using masked prediction on audio signals in order to extract representations on the audio for each frame. These SSL techniques typically extract representations at a fixed frame rate at around 50 Hz, which is fairly finegrained and suggests that these representations are highly correlated with sub-phonemic structures (Hsu et al., 2021; Cho et al., 2023; Abdullah et al., 2023; Sicherman & Adi, 2023; Baevski et al., 2021).
Speech tokenization Clustering and/or quantizing these SSL representations can provide speech tokens that are used for acoustic unit discovery (Hallap et al., 2022), speech recognition (Baevski et al., 2021; Chang et al., 2024), speech synthesis (Polyak et al., 2021; Hassid et al., 2024), language modeling (Lakhotia et al., 2021; Borsos et al., 2022; Hassid et al., 2024; Zhang et al., 2023), and translation (Lee et al., 2022; Li et al., 2023). However, these tokens severely suffer from high sampling rates, which makes the downstream models hard to scale, and struggle to learn long-range dependencies and higher-level linguistic structures due to the lack of explicit word boundaries and longer sequences. These caveats can be greatly improved by tokenizing speech at syllable-level granularity. [^3] In English, the typical speaking rate is 4-5 syllables per second.
Syllabic structure in speech SSL Previous studies have demonstrated that syllabic structure can be induced by SSL (Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024). Peng et al. (2023) shows that syllabic structure in SSL features can be induced by jointly training with images and spoken captions. SDHuBERT (Cho et al., 2024b) demonstrates that such syllabic induction can be free of other modalities, with a sentence-level SSL. Komatsu & Shinozaki (2024) combined frame-wise distillation with speaker augmentation in order to derive syllabic segments. All these methods then utilize an agglomeration algorithm on top of the learned features to infer syllable boundaries, and extract syllable embeddings by averaging frames within detected segments. However, all of these prior studies induce syllabic structures through indirect ways, resulting in noisy syllable boundaries. Moreover, it is unclear whether the discovered syllables are valid speech representations or tokens. Our approach greatly improves the quality of the segments. Moreover, we demonstrate the efficacy and validity of syllabic tokens through experiments.
@@ 3.1 Self-Segmentation Distillation
Figure 1: Overview of self-segmentation distillation. Sylber is trained with frame-wise regression on pseudo segment targets, obtained by an unsupervised segmentation algorithm on the teacher outputs.
Sylber is trained by a novel SSL framework, self-segmentation distillation, that imposes more explicit inductive bias of segment structure in feature representations by directly solving the speech segmentation problem. The diagram of our training process is depicted in Figure 1. Specifically, we use SDHuBERT (Cho et al., 2024b) as a starting point, and leverage its unsupervised syllable segments as pseudo targets of segmentation. The target segment labels are continuous embeddings averaged across frames within each segment that are found by an unsupervised segmentation algorithm. We use self-supervised knowledge distillation, where the teacher is an exponential moving average (EMA) of the student model (Grill et al., 2020; Caron et al., 2021; He et al., 2020). The target segment and segment embedding are extracted from the teacher, making the learning process self-supervised and free of labels. The loss objective is a frame-wise regression loss that minimizes the Mean Squared Error (MSE) between the output features at each frame and the target from the corresponding segment. Non-speech frames are regressed to zero. See Appendix A.1.1 for a formal definition.
Other than EMA, this learning objective is free of techniques that prevent collapse (e.g., contrastive learning, target recentering, masked prediction, etc). However, we find that initializing Sylber weights with SDHuBERT can avoid collapse even though a naive regression is highly vulnerable to collapse. Additionally, we include a denoising objective similar to Chen et al. (2022) to improve robustness of the model, where 20% of the batch inputs for the student are mixed with environmental noise (Reddy et al., 2021) or other speech audio. This additional denoising is not a primary source of learning as a syllabic structure is readily visible without it (Appendix A.1.7).
@@ 3.2 Linear time Greedy segmentation algorithm
Figure 2: Frame-wise similarity matrix of raw features measured by dot product. For HuBERT and SDHuBERT, features from the ninth Transformer layer are extracted. As we can see, Sylber shows extremely salient syllabic structure that is aligned with the ground truth syllable boundaries.
The result of our self-segmentation distillation induces a framewise speech representation that exhibits a segmented structure as seen in the frame-wise similarity matrix (Figure 2, right). As we can see, our method produces a clean and robust segment structure that we can take advantage of to design a linear-time, greedy audio segmentation algorithm (also shown in Algorithm 1).
The algorithm involves three linear passes through the audio embeddings. The first step thresholds all of the embeddings by their L2 norm. This step allows us to differentiate between speech and non-speech segments. The second step is a monotonic agglomeration process where we sweep through each embedding and aggregate them into segments. Adjacent frames are merged together into a segment as long as their cosine similarity goes above a predefined merge threshold. This can be done in single pass without constructing the entire similarity matrix by greedily creating a new segment once a frame with a similarity below the threshold is seen.
The greedy segmentation algorithm can sometimes make some errors by shifting some frames, so a third pass is used to refine the boundaries of adjacent segments. For each boundary, a local search range is defined from the midpoint of the previous segment to the midpoint of the sub-sequence segment. From here, we can compute the cosine similarity between each frame and the averages of the two segments. For each candidate boundary in the search range, we compute an aggregate cosine similarity score between each frame and the assigned segment and maximize this sum to find the optimal boundary between adjacent segments.
@@
Each one of these steps can be implemented with $O(n)$ complexity, so the entire segmentation algorithm has linear complexity with respect to the audio sequence length. As we can see in Table 1, this is significantly more efficient that previous segmentation approaches (Peng et al. (2023); Cho et al. (2024b); Komatsu & Shinozaki (2024)) which all have $O(n^{2})$ complexity.
4 Evaluating Categorical Perception in Speech Representation
Figure 3: A. Overview of articulatory interpolation of rhyming words when interpolating $\alpha\in[0,1]$. B. Hypothetical curves of categorical (solid lines) and non-categorical (dashed lines) embeddings. C. Similarity curves examples from Melspectrogram (Mel), HuBERT layer 9, and Sylber. Sylber consistently shows highly categorical perception, drawing a sharp boundary in continuum between words.
Previous SSL-based tokens suffer from high redundancy in token vocabulary (Sicherman & Adi, 2023). This indicates that the SSL features densely tile the phonetic space without a clear boundary, resulting fine-grained, sub-phonemic units when clustered. Thus, to be better tokenized, the features should have distinct boundaries in their embedding space. This is well-aligned with categorical perception, a linguistic theory, which argues that human speech perception draws a categorical boundary in a continuum of speech sounds (Liberman et al., 1957; Pisoni & Lazarus, 1974; Harnad, 2003).
Inspired from this linguistic theory, we simulate interpolation between two rhyming words to probe the embeddings of speech SSL models to check whether they are categorical. Specifically, monosyllabic words are recruited where a single consonant is different at the front or back of the syllable (onset or coda). We make the contrast to be switching one of phonological properties: nasality, voicedness, or place (e.g., “b”all vs “m”all, “d”own vs ”t”own, or “l”est vs “r”est, respectively). We do not include vowel contrasts since categorical perception of vowels is not as consistent as consonants (Pisoni, 1973; Pisoni & Lazarus, 1974). We consider 13 types of such difference and simulate 4 pairs for each type, resulting 52 word pairs in total. The details of the difference types and the full list of word pairs can be found in Appendix A.4. To simulate a continuum between words, we utilize Articulatory Encodec (Cho et al., 2024c) which allows direct editing in the physical articulatory space (Figure 3-A). We first generate audio using the an off-the-shelf TTS API. [^4] We use the TTS service in Vertex AI (https://cloud.google.com/vertex-ai) with a default female voice. We extract articulatory features from the speech, which are then temporally aligned by dynamic time warping to either end if necessary. We sample 51 equidistant samples in the linear interpolation between two words, where each end is manually adjusted to make the perceptual boundary drawn approximately in the middle ($\alpha=0.5$), which can be heard here. The pitch and loudness are also controlled to be at the same level. More details about Articulatory Encodec can be found in Section A.1.3.
Given a speech representation model, we extract features for each interpolating point between words in each pair. We calculate the similarity between interpolating features with features from either end, forming a likelihood curve along the interpolation. Hypothetically, if the representation is categorical, the likelihood curves should show a sharp transition at the boundary (Figure 3-B). If the embeddings are not categorical and tracing the interpolation, the curves would show “X” pattern as the dashed line in Figure 3-B. To quantify discriminability, we measure an empirical risk of wrong discrimination. Given words at the left and right ends in interpolation, $x_{L},x_{R}\in W$, the probability of being the left word given interpolating factor $\alpha$ is defined as $p(x_{L}|x_{\alpha})=\frac{\text{sim}(x_{L},x_{\alpha})-\text{offset}_{L}}{\text{sim}(x_{L},x_{\alpha})-\text{offset}_{L}+\text{sim}(x_{R},x_{\alpha})-\text{offset}_{R}}$. The probability of being the right word is symmetrically defined. We need to subtract an offset due to the high base similarity, as the words are rhyming pairs, such that $\text{offset}_{A}=\text{min}_{\alpha\in[0,1]}\text{sim}(x_{A},x_{\alpha})$. Then the empirical risk can be defined as $L_{\text{Disc}}(q|x_{L},x_{R}):=\mathbb{E}_{\alpha\in[0,1]}\mathbbm{1}_{\alpha<q}p(x_{R}|x_{\alpha})+\mathbbm{1}_{\alpha\geq q}p(x_{L}|x_{\alpha})$, for the decision boundary at $q\in[0,1]$. The optimal boundary can be drawn by minimizing the risk, $\alpha^{*}=\text{argmin}_{q\in[0,1]}L_{\text{Disc}}(q|x_{L},x_{R})$. Discriminability index (DI) is then defined as the risk at the optimal boundary, averaged over word pairs:
If the embeddings are categorical, DI will be close to 0. If they are non-categorical with X-shaped curves, DI will be 0.25. The maximum value of DI is 0.5, which would be random chance discrimination. This will be discussed later in §6.4 in detail.
@@ 5.1 Experimental Setup
Tokenization To tokenize speech, we apply the aforementioned segmentation algorithm (Section 3.2) to get unsupervised speech segments. The features within segments are averaged to form continuous speech tokens at a syllable granularity (4-5 syllables per second). We apply a simple k-means clustering on the features with several different vocab sizes (5K, 10K, and 20K). These cluster sizes are larger than what is used by other SSL-based clustering techniques (usually around 50-500 clusters), which is necessary since our features are more closer to syllables than phonemes; similar to how vocabulary sizes for BPE based tokenizers are significantly larger than the number of characters. However, these syllabic tokens have a significantly lower temporal resolution compared to previous SSL-based tokens, which leads to improvements to efficiency (see Section 6.2).
Token-to-speech If our syllabic tokens are valid speech tokens, we should be able to reconstruct intelligible speech from them. We train a Conditional Flow-matching (CFM) (Lipman et al., 2022; Le et al., 2024) model to generate low-level speech features that can be converted to speech audio. We utilize Articulatory Encodec (Cho et al., 2024c), an encoding-decoding framework that encodes speech into articulatory features and speaker identity, and decodes them back to speech waveform. The articulatory features are composed of spatiotemporal displacements of vocal tract articulators and source features that represent loudness and pitch. Cho et al. (2024c) empirically prove that these articulatory features are speaker agnostic provided that pitch is normalized, while allowing full-reconstruction to speech. We normalize pitch by dividing by speaker’s mean pitch and logscaling, following (Kharitonov et al., 2021). Since SSL-based speech tokens generally lack speaker information (Polyak et al., 2021; Wang et al., 2023), we aim to reconstruct these speaker-agnostic articulatory features from the syllabic tokens. The token embeddings are restored by the k-means codebooks, and expanded to the durations of the original segments. The non-speech frames are filled with zeros. For the case without quantization, the segment-averaged features are used. At inference time, we input the generated articulatory features with speaker embeddings extracted from original speech to the decoder of Articulatory Encodec to generate audio. See Appendix A.1.3 for the implementation and training details.
Unit LM Following Lakhotia et al. (2021), we train an autoregressive unit language model (uLM) using the syllabic tokens. The model has the same architecture as GSLM (Lakhotia et al., 2021), which is a decoder-only Transformer with 12 layers.
Datasets LibriSpeech (Panayotov et al., 2015) is used for pretraining Sylber, k-means clustering, and for training the uLMs. LibriTTS-R (Koizumi et al., 2023) is used for training the CFM models. For training Articulatory Encodec, we use an extended dataset that includes LibriTTS-R, LibriTTS (Zen et al., 2019), and EXPRESSO (Armougom et al., 2006).

### v2

@@ Abstract
Syllables are compositional units of spoken language that efficiently structure human speech perception and production. However, current neural speech representations lack such structure, resulting in dense token sequences that are costly to process. To bridge this gap, we propose a new model, Sylber, that produces speech representations with clean and robust syllabic structure. Specifically, we propose a self-supervised learning (SSL) framework that bootstraps syllabic embeddings by distilling from its own initial unsupervised syllabic segmentation. This results in a highly structured representation of speech features, offering three key benefits: 1) a fast, linear-time syllable segmentation algorithm, 2) efficient syllabic tokenization with an average of 4.27 tokens per second, and 3) novel phonological units suited for efficient spoken language modeling. Our proposed segmentation method is highly robust and generalizes to out-of-domain data and unseen languages without any tuning. By training token-to-speech generative models, fully intelligible speech can be reconstructed from Sylber tokens with a significantly lower bitrate than baseline SSL tokens. This suggests that our model effectively compresses speech into a compact sequence of tokens with minimal information loss. Lastly, we demonstrate that categorical perception—a linguistic phenomenon in speech perception—emerges naturally in Sylber, making the embedding space more categorical and sparse than previous speech features and thus supporting the high efficiency of our tokenization. Together, we present a novel SSL approach for representing speech as syllables, with significant potential for efficient speech tokenization and spoken language modeling.
@@ 1 Introduction
A major bottleneck of the inefficiency in modeling spoken language is a lack of structure in current neural speech representations. Unlike text, there is no clear delimiter nor orthographic symbol in speech audio, which are crucial in efficient and scalable processing as evidenced in the text domain. However, human speech perception is structured as being segmented (Greenberg, 1998; Oganian & Chang, 2019; Gong et al., 2023) and categorical (Liberman et al., 1957; Pisoni, 1973; Pisoni & Lazarus, 1974). We argue that the machine representation of speech should resemble these cognitive structures to allow similar efficiency as text processing. A natural segmented structure of speech is a syllable, which organizes speech sounds in time (MacNeilage, 1998; Greenberg, 1998), and ideally, the embedding of a syllable should represent contents in a categorical way to be symbolized efficiently.
To this end, we propose a novel SSL framework that induces clean and robust syllabic structures in speech representations. Specifically, we build on top of a previous self-supervised syllable learning model, SDHuBERT (Cho et al., 2024b), and iteratively refine the syllabic segments that naturally arise from the model. Unlike the original model, which induces syllable structure as a byproduct of sentence-level SSL, we directly impose syllabic structures by regressing features against unsupervised syllable segments extracted from a teacher model which is initially set as the training model. We call the resulting model Sylber (Syllabic embedding representation).  [^1] The code is available here: https://github.com/Berkeley-Speech-Group/sylber.
The features from Sylber exhibit salient syllabic structure—showing a flat, consistent output within each segment and distinctive from other syllables (Figure 2, right). This enables a fast, linear time algorithm for segmenting these features. Moreover, this allows more accurate boundary detection and clustering that is more coherent with ground truth syllables than previous approaches. Syllabic tokens quantized from Sylber features show significantly lower frequency at an average of 4.27 token/second (Tok/s), and can be used to synthesize fully intelligible speech. [^2] Audio samples are at https://berkeley-speech-group.github.io/sylber. Furthermore, spoken language models based on syllabic tokens show comparable or better performance than the baselines with a similar resource setting, in learning lexicons and syntax.
@@
  • We propose Sylber, a novel SSL framework that imposes salient and robust syllabic structure in speech representation.
  • Sylber outperforms previous approaches in syllable detection and discovery with a more efficient segmentation algorithm with $O(n)$ time complexity.
  • The syllable segmentation by Sylber is generalizable to noisy conversational speech and even to unseen languages (Spanish and Mandarin) while being trained only on English audiobook
  • We use this model to build a new dynamic speech tokenization scheme that has significantly lower sampling rate as 4.27 Tok/s on average, 6-7 times improvement over HuBERT tokens.
  • We demonstrate that fully intelligible speech can be reconstructed from syllabic tokens, and that these units are suited for lexical and syntactic understanding.
@@ 2 Related Work
Self-supervised learning in speech Self-supervised learning (SSL) has been leveraged to learn rich representations from large, unlabeled speech corpora (Hsu et al., 2021; Baevski et al., 2020; Chen et al., 2022; Chung et al., 2021; Mohamed et al., 2022). Notably, HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022) are pretrained using masked prediction on audio signals in order to extract representations on the audio for each frame. These SSL techniques typically extract representations at a fixed frame rate at 50 Hz, which is fairly finegrained and suggests that these representations are highly correlated with sub-phonemic structures (Hsu et al., 2021; Cho et al., 2023; Abdullah et al., 2023; Sicherman & Adi, 2023; Baevski et al., 2021).
Speech tokenization Clustering and/or quantizing these SSL representations can provide speech tokens that are used for acoustic unit discovery (Hallap et al., 2022), speech recognition (Baevski et al., 2021; Chang et al., 2024), speech synthesis (Polyak et al., 2021; Hassid et al., 2024), language modeling (Lakhotia et al., 2021; Borsos et al., 2022; Hassid et al., 2024; Zhang et al., 2023), and translation (Lee et al., 2022; Li et al., 2023). However, these tokens severely suffer from high sampling rates, which makes the downstream models hard to scale, and struggle to learn long-range dependencies and higher-level linguistic structures due to the lack of explicit word boundaries and longer sequences. These caveats can be greatly improved by tokenizing speech at syllable-level granularity, yielding far shorter token lengths than the existing methods. For example, in English, the typical speaking rate is 4-6 syllables per second.
Syllabic structure in speech SSL Previous studies have demonstrated that syllabic structure can be induced by SSL (Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024). Peng et al. (2023) shows that syllabic structure in SSL features can be induced by jointly training with images and spoken captions. SDHuBERT (Cho et al., 2024b) demonstrates that such syllabic induction can be free of other modalities, with a sentence-level SSL. Komatsu & Shinozaki (2024) combined frame-wise distillation with speaker augmentation to derive syllabic segments. All these methods then utilize an agglomeration algorithm on top of the learned features to infer syllable boundaries, and extract syllable embeddings by averaging frames within detected segments. However, these previous studies induce syllabic structures in indirect ways, resulting in noisy syllable boundaries. Moreover, it is unclear whether the discovered syllables are valid speech representations or tokens. Here, our approach greatly improves the quality of the segments, and we first demonstrate the efficacy and validity of syllabic tokens through various experiments.
@@ 3.1 Self-Segmentation Distillation
Figure 1: Self-segmentation distillation.
Sylber is trained by a novel SSL framework, self-segmentation distillation, that imposes more explicit inductive bias of segment structure in feature representations by directly solving the speech segmentation problem (Figure 1). Specifically, we use SDHuBERT (Cho et al., 2024b) as a starting point, and leverage its unsupervised syllable segments as pseudo labels for segmentation. The target segment labels are continuous embeddings averaged across frames within each segment which are extracted from a teacher model. The teacher model is set as the initial student model and fixed during training, where we have two training stages—use the initial segments by SDHuBERT and its model weights; and update the teacher model and segments using the student model trained in the first stage.
The loss objective is a frame-wise regression loss that minimizes the Mean Squared Error (MSE) between the output features at each frame and the target embeddings from the corresponding segment. Non-speech frames are regressed to zero, which are marked by norm thresholding by SDHuBERT (Cho et al., 2024b), and segments with low waveform amplitude (Appendix A.1.7). A formal definition of this distillation loss is described in Appendix A.1.1.
In addition to the distillation loss, we include a denoising objective similar to Chen et al. (2022) to improve robustness of the model, where 20% of the batch inputs for the student are mixed with environmental noise (Reddy et al., 2021) or other speech audio (Appendix A.1.3). This additional denoising is not a primary source of learning as a syllabic structure is readily visible without it, which is qualitatively shown in Appendix A.2.2. Furthermore, the model training is not sensitive to the choice of hyperparameters or model initialization (more discussed in Appendix A.2.6).
@@ 3.2 Linear time Greedy segmentation algorithm
Figure 2: Frame-wise similarity matrix of raw features measured by dot product. For HuBERT and SDHuBERT, features from the ninth Transformer layer are extracted. Sylber shows extremely salient syllabic structure that is aligned with the ground truth syllable boundaries, with clear null activations in non-speech frames.
The result of our self-segmentation distillation induces a framewise speech representation that exhibits a segmented structure as seen in the frame-wise similarity matrix (Figure 2, right). As we can see, our method produces a clean and robust segment structure that we can take advantage of to design a linear-time, greedy audio segmentation algorithm. The algorithm involves a monotonic agglomeration process where we sweep through each embedding and aggregate them into segments. While sweeping, the adjacent frames are merged together into a segment as long as their cosine similarity goes above a predefined merge threshold. The frames with L2 norm lower than a norm threshold are regarded as non-speech, and terminate agglomerating whenever encountered while sweeping. The greedy segmentation algorithm can sometimes make some errors by shifting some frames. Therefore, an additional pass is used to locally refine the boundaries by maximizing similarities between frames and assigned segments. See Appendix A.1.2 for details.
@@
Each one of these steps can be implemented with $O(n)$ complexity, so the entire segmentation algorithm has linear complexity with respect to the audio sequence length. This allows fast online segmentation and is significantly more efficient than previous segmentation approaches (Peng et al. (2023); Cho et al. (2024b); Komatsu & Shinozaki (2024)) which all have $O(n^{2})$ complexity as shown in Table 1. In Section 5.1, we show that this algorithm is on-par with previous $O(n^{2})$ algorithms when applied to Sylber and other ablations involving previous models.
@@ 4.1 Experimental Setup
Tokenization To tokenize speech, we apply the aforementioned segmentation algorithm (Section 3.2) to get unsupervised speech segments. The features within segments are averaged to form continuous speech tokens at a syllable granularity (4-5 syllables per second). We apply a simple k-means clustering on the features with different vocab sizes (5K, 10K, and 20K). These cluster sizes are larger than what is used by other SSL-based clustering techniques (usually around 50-2K clusters), which is necessary since our features are closer to syllables than phonemes; similar to how vocabulary sizes for BPE based tokenizers are significantly larger than the number of characters. However, these syllabic tokens have a significantly lower temporal resolution compared to previous SSL-based tokens, which leads to improvements in efficiency (see Section 5.2).
Token-to-speech If our syllabic tokens are valid speech tokens, we should be able to reconstruct intelligible speech from them. We train a Conditional Flow-matching (CFM) (Lipman et al., 2022; Le et al., 2024) model to generate interpretable articulatory features that can be converted to speech audio using SPARC (Cho et al., 2024c). These articulatory features are speaker agnostic provided that pitch is normalized, while allowing full-reconstruction to speech. Since SSL-based speech tokens generally lack speaker information (Polyak et al., 2021; Wang et al., 2023), we aim to reconstruct these speaker-agnostic articulatory features from the syllabic tokens. See Appendix A.1.4 for the implementation and training details.
Unit LM Following Lakhotia et al. (2021), we train an autoregressive unit language model (uLM) using the syllabic tokens. The model has the same architecture as GSLM (Lakhotia et al., 2021), which is a decoder-only Transformer with 12 layers. Additional details are in Appendix A.1.6.
Datasets LibriSpeech (Panayotov et al., 2015) is used for training Sylber, and k-means clustering. For training the uLMs, we use either LibriSpeech or LibriLight (Kahn et al., 2020), and separately report performance. LibriTTS-R (Koizumi et al., 2023) is used for training the CFM models.